40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 176 of 344  Not logged in ELOG logo
ID Date Author Type Category Subject
  8492   Thu Apr 25 17:56:28 2013 RijuConfiguration PD frequency response

 [Eric, Riju]

Today we have routed the fibers from 1x16 fiber splitter to POX table for POX11 PD and POP55 PD. Also we labeled the fibers on AP table, they have been fixed on the table. The photo of the table after work is attached here. We will do it for POX table tomorrow. 

  8491   Thu Apr 25 10:19:10 2013 KojiSummaryLSCLocking activity on Apr 24th

Last night I worked on the several locking configurations:

General preparations / AS table inspection

- The AS beam looked clipped. I went to the AP table and confirmed this is a clipping in the chamber.
  This may be fixed by the invacuum PZTs.

Modulation frequency tuning

RFPD Mon of the MC demodulator was check with the RF analyzer. Minimized the 25.8MHz (=55.3-29.5MHz) peak by changing the marconi freq.
This changed the modulation freq from 11.066147MHz to 11.066134MHz. This corresponds to the change of the MC round-trip length from
27.090952m to 27.090984m (32um longer).

Michelson tests

- I wonder why I could not see good Michelson signal at REFL ports.

- I roughly aligned the Michelson. On the AP table, the RF analyzer was connected to the REFL11 RF output.
  By using "MAX HOLD" function of the analyzer, I determined that the maximum output of the 11.07MHz peak
  was -61.5dBm.

- I went to the demodboard rack. I injected -61dBm from DS345 into the RFEL11 demodboard. This produced
  clean sinusoidal wave with the amplitude of 4 count. The whitening gain was 0dB.

- The output from the PD cable was -64.0dBm. So there is ~2.5dB loss in the cable. Despite this noise, the demodulation
  system should be sufficiently low noise. i.e. the issue is optical

- The Michelson was locked with AS55Q. And the REFL11 error signals were checked.Fringe like feature was there.
  This suggested the scattering from the misaligned PRM. The PRM was further misaligned. Then some reasonable
  (yet still noisy) Michelson signal appeared. (Usual misaligned PRM is not at the right place)

  Q. How much scattering noise (spurious cavity between PRM and the input optics) do we have when the PRM is aligned?
  Q. Where should we put the glass beam dumps in the input optics?
  Q. Can we prepare "safe" misaligned place for the PRM with the beam dump?

- The Michelson was locked with REFL11Q. From the transfer function measurement, the gain difference between AS55Q (whitening gain 24dB)
  and REFL11Q was 32dB. The whitening gain was 0dB. In fact I could not lock the Michelson with the whitening gain 33dB (saturation???)
  The element in the Input matrix was 1, The gain of the servo was +100. BS was actuated.

Coupled cavity tests

- At least REFL11 is producing reasonable signals. So what about the other REFL ports? The Michelson signals in the other frequencies
  were invisible. So I decided to use three-mirror coupled cavity with the loss PRC.

- Aligned X arm, Misaligned ETMX, ITMY. Aligned PRM.

- Locked the PRM-ITMX cavity with REFL11 and REFL33.

- Aligned ETMX. If I use REFL11I for the PRC locking, I could not lock the coupled cavity. But I could with REFL33I.
  This is somewhat familiar to me as this is the usual feature of the 3f signal.

- The coupled cavity could be locked "forever". To realize this I needed to tweak the normalization factor from 1.0 to 1.6.
  Q. How does the coupled cavity change the response of the cavity? Can we compensate it by something?
  Q. Measure open loop transfer functions to check if there is any issue in the servo shapes.

- Transmission during the lock is 3.2 while the nominal TRX with PRM misaligned was 0.93.
  This corresponds to power recycling gain of 0.17.

 - X arm:

    - Source: POX11I, phase 79.5 deg, whitening gain 36dB
    - Input matrix: POX11I->1.0->XARM, Normalization TRX*1.60
    - XARM servo gain +0.8, actuation ETMX
    - XARM trigger 0.25 up, 0.05 down. XARM Filter trigger untouched.

- PRC: (sideband locking)
    - Source: REFL33I, phase -34.05 deg, whitening gain 30dB
    - Input matrix: REFL33I->1.0->PRCL, Normalization None
    - PRCL servo gain +4.0, actuation PRM
    - PRCL trigger None

- Same test for the Y arm. At the moment ETMY did not have the OPLEV.
  Same level of transmission (~3.3)

 - Y arm:

    - Source: POY11I, phase -61.00 deg, whitening gain 36dB
    - Input matrix: POY11I->1.0->YARM, Normalization TRX*2.1
    - YARM servo gain +0.25, actuation ETMX
    - YARM trigger 0.25 up, 0.05 down. YARM Filter trigger untouched.

- PRC: (sideband locking)
    - same as above

Sideband PRMI attempt

    - Now I got some kind of confidence on the REFL33 signal.
    - So I tried to get any stable setup for sb PRMI, then to find any reasonable MICH signals anywhere else than AS55Q.
    - With REFL33I(PRCL) & AS55Q(MICH), I got maximum ~10sec lock. It regularly locked. It was enough long to check
      the spectrum on DTT. But it was not enough long to find anything about the MICH signals at the REFL ports.

    - I tried REFL33Q for MICH. The lock was even shorter but could lock for 1~2 sec.

    Q. What is the cause of the lock loss? I did not see too much angluar fluctuation. The actuation was also quiet (below 10000).

- PRCL: (sideband locking)
    - Same as above except for
      - the PRCL servo gain +0.05, No limitter at the servo output.
      - Trigger POP22I (low pass filtered by LP10) 20 up, 3 down

- MICH:
    - AS55Q -24.125 24dB -> x1.0 -> MICH -0.7, No limitter -> ITMX/Y differential
    or
    - REFL33Q -34.05dB -> x2.0 -> MICH same as above
    - For both case, trigger POP22I (low pass filtered by LP10) 20 up, 3 down

 

At this point Jenne came back from dinner. Explained what I did and handed over the IFO.

  8490   Thu Apr 25 04:10:09 2013 JenneUpdateLockingMICH_CTRL drifting away??

Koji is elogging separately of his exploration of different configurations.  The lock stretch that I'm looking at here uses the same parameters as Koji had for PRMI sb lock, using AS55Q for MICH and REFL33I for PRCL, with MICH gain of -0.8 and PRCL gain of 0.05 .

All of these plots are the same few second lock stretch, with different zooming.  Jamie's super-sweet getdata python script only accepts integers for the start time and duration parameters, so lots of this zooming happened by hand, but I tried to always keep the time axis aligned within each screenshot.  Sometimes the plot axis labels say differently, but they're lying to you.

Plot 1:  gps start time is 1050915916, duration = 6 seconds.  Overall view of the lock stretch.

1050915916-6.png

Plot 2:  gps start time is 1050915921, duration = 1 second.  We're looking at the lockloss that happens at the left side of the plots.

1050915921-1.png

Plot 3:  zoomed in (along the time-axis) version of plot 2, so much shorter time duration.  Some zooming on y-axes.

1050915921-zoom.png

Plot 4:  zoomed in (along y-axes) version of plot 2.

1050915921-1-zoom.png

It seems to me from these plots that maybe MICH CTRL is drifting away?  It seems like we lose the MICH lock, and that destroys the whole thing. 

Koji made some comments to me earlier, regarding his work this evening, that the MICH signal quality is poor in general, and that we should calculate/think about changing our schnupp asymmetry. 

  8489   Thu Apr 25 03:35:28 2013 JenneUpdateLockingAngular motion does not explain RIN

Den made a nice elog about the PRMI RIN that we see a few weeks ago:  8464.  The RIN that we're seeing is typically about ~30%.  The question at hand is: what is causing this power fluctuation, and more specifically, is it the angular motion of the mirrors?

I find that no, the angular motion that we see does not explain the RIN that we see.

In the attached Mathematica notebook, I calculate the power lost due to angular misalignments of one or more mirrors.  (Math comes from Appendix A of Keita's thesis.)

From calibrated oplev spectra, our mirrors are moving about 1 microradian (RMS, which is dominated by low frequencies).  From a super sophisticated "draw on the TV, then measure" method (details below), I have estimated that the maximum static misalignment that we're seeing is about 2 microradians.

With all of this, I find that for a g-parameter of 0.94, the power lost due to misalignments should, at maximum, be 0.6%.  I need a g-parameter of 0.995 to get a power loss of 23%.  Alternatively, if I take the derivative of the power coupling function, to find the static misalignment at the steepest slope of the curve (and thus, the place where any AC misalignment would have the most effect), for 1urad of AC misalignment, I get 40% power loss. 

So, in order for the AC angular motion that we see to explain the RIN that we see, either our mirrors are very, very misaligned (so much so that we couldn't really be locking), or our cavity is much closer to unstable than expected from Jamie's calculations.  Since both of these cases (static misalignment or incorrect g-parameter calculation) have to be taken to extremes before they approximate the RIN that we see, I do not think that this power loss is due to angular fluctuations.

This means that we have to think of another potential cause for this RIN that we're seeing.

Details on the "draw on TV and measure" technique for determining static cavity misalignments:  Looking at the POP camera view, with the PRM significantly misaligned, I traced the straight-through beam spot.  I then restored the PRM, and during several momentary locks, I traced the beam spot, which I took to be the saturated area of the camera.  The idea here is that the straight-through beam represents the incident beam axis, while the locked beam represents the cavity axis.  I'm assuming that the camera image plane is at the face of PR2. I approximately found the center of each of my tracings, and found them to be ~1/4 inch apart.  I also measured the "spot size" of the sideband-locked PRMI, and found it to be ~3.5 inches.  So, very roughly, the ratio of (distance between spots)/(size of beam) is ~0.07. This corresponds to a static misalignment of either the ITM or the PRM of ~2urad, rounding up. (I use the Jamie's calculated g-parameters from elog 8316, the case of flipped PR2, tangential = 0.94 to calculate the effective RoC of the PRM). 

  8488   Thu Apr 25 00:59:37 2013 RijuUpdateRF SystemPD frequency response

Quote:

I think you have the splitter that splits the RF signal from the network analyzer in the wrong place. 

Usually you split the signal immediately after the RF Out, so that half of the signal goes to the A-input of the Analyzer, and the other half goes to your controller (here, the laser diode controller).  Then you would take the output of your controller and go straight to the actual laser diode, with no splitting in this path.

 Here our device under test is the photodiode. So for the reference I wanted to retain the response of the laser diode controller. Otherwise I have to consider the transfer function of that LDC too. I may check both the options at the time of experiment.

Thanks

  8487   Wed Apr 24 18:51:12 2013 KojiConfigurationoptical tablesPD frequency response

The fibers should be routed beneath the electrical cables.
They should be fixed on the table for strain relieving.
The slack of the fibers should be nicely rolled and put together at the splitter side.

These are expected to be done next time when the fiber team work around the table.

We also expect to have the table photo every time the work of the day is finished.

  8486   Wed Apr 24 15:27:59 2013 ManasaUpdate40m UpgradingEndtable upgrade for auxiliary green laser : progress: New layout

Layout that will be improved upon over the next few days.

Things that need to be updated:
1. Waist size at all optics
2. Beam size at detectors and choice of lenses
3. IPANG & green PD proposed positions

 ETMY_endtable_New.png

  8485   Wed Apr 24 14:36:06 2013 JenneUpdateRF SystemPD frequency response

I think you have the splitter that splits the RF signal from the network analyzer in the wrong place. 

Usually you split the signal immediately after the RF Out, so that half of the signal goes to the A-input of the Analyzer, and the other half goes to your controller (here, the laser diode controller).  Then you would take the output of your controller and go straight to the actual laser diode, with no splitting in this path.

  8484   Wed Apr 24 14:24:40 2013 RijuUpdate PD frequency response

 Here I am attaching the first schematic diagram of the PD frequency response set-up, I will keep updating it with relevant informations with the progress of the work.

Description: Our objective is to set-up one simultaneous transfer-function measurement system for all the RF-PDs present in 40m lab. A diode laser will be used to illuminate the PDs. The diode laser output will be divided by 1x16 fiber splitter and will be sent to all the PDs through single-mode fiber. The transfer function of the PDs will be measured using network analyzer(Agilent 4395A). The output of the PDs will be fed to network analyzer via one RF-switch. The diode laser will be controlled by the controller ILX LDC 3744C. The scanning frequency signal will be fed to this controller from network analyzer through its external modulation port. The output of the controller will be splitted  into two parts: one will go to laser diode and the other will be used as reference signal for network analyzer.

 

 

  8483   Wed Apr 24 14:20:49 2013 KojiUpdateCDSFE Web view not updated?

The FE web view seems not up-to-date, does it? ( maybe for a year)

https://nodus.ligo.caltech.edu:30889/FE/c1mcs_slwebview_files/index.html

  8482   Wed Apr 24 00:44:33 2013 KojiUpdateSUSPRMI locked, ITMX pitch OpLev ringing up

I tried to reproduce the locking situation described in this entry tonight.
The momentary lock was regularly seen but there was no stable lock.

I wonder why the actuators are always saturated. The feedback signals have the dominant component at ~400Hz.

It would also be nice if the servos have some immunity to gain fluctuation.

I didn't check how the situation of the AP table is. I'll look into some details tomorrow.

  8481   Wed Apr 24 00:42:07 2013 JenneUpdateSUSPRMI locked, ITMX pitch OpLev ringing up

Koji is working on PRMI locking, and while he was doing that I glanced at the oplevs' spectra for the ITMs and PRM.

I found that when the PRMI was locked (for only 1 second or so max lock time) on the 55MHz sideband, and the error signals show a big peak around 400Hz (definitely audible in the control room), the only OpLev that I see a similar peak in is ITMX pitch. 

In the plot below, I have grabbed a time when the PRMI was flashing as the black reference traces, and then a time when the PRMI was locked as the active traces.  You can see that there is a similar peak in both REFL55I and ITMX_OL_PIT when the cavity is locked. 

PRMI_locked_ITMXpitOpLevRingingUp.pdf

  8480   Tue Apr 23 22:59:05 2013 ranaConfigurationSUSOptical Lever Gains normalized

Due to the recent addition of cal factors in the OL error points, the OLPIT_GAIN and OLYAW_GAIN have been reduce to tiny numbers (e.g. 0.002).

Since our MEDM only shows 3 digits past the decimal point by default, it makes more sense to have the gains around 1.

So I reduced the gains in all of the FM1 filters from 1000 to 1 and multiplied the GAIN values by 1000 (using ezcastep) to compensate.

All of the active optics seem to be behaving as before. Haven't tested ETMs or SRM yet.

  8479   Tue Apr 23 22:10:54 2013 ranaUpdateComputersNancy

controls@rosalba:/users/rana/docs 0$ svn resolve --accept working nancy
Resolved conflicted state of 'nancy'

  8478   Tue Apr 23 16:31:13 2013 EricConfiguration PD frequency response

[Eric, Riju]

Summary: Routing Fibers on AP table for Photo Diode Frequency Response Measurement System

Objective: We are to set-up one simultaneous transfer-function measurement system for all the RF-PDs present in 40m lab. A diode laser output is to be divided by 1x16 fiber splitter and to be sent to all the PDs through single-mode fiber. The transfer function of the PDs will be measured using network analyzer. The output of the PDs will be fed to network analyzer via one RF-switch.

Work Done So Far: We routed the fibers on AP table. Fibers from RF PDS - namely  MC REFL PD, AS55, REFL11, REFL33, REFL55, REFL165, have been connected to the 1x16 fiber splitter. All the cables are lying on the table now, so they are not blocking any beam.

We will soon upload the schematic diagram of the set up.

 

Missing Component: Digital Fiber Power Meter, Thorlab PM20C

 

 

  8477   Tue Apr 23 16:17:45 2013 DenUpdatePEMseismometer isolation kit in place

Quote:

 

 The carpenter shop finished the installation of the 3 granite bases.Rapid Set Cement All high strength non-shrink grout was used.

 Compressive strength  3000 PSI at 1 hour and 9000 PSI at day 28 The janitor is still cleaning up after them at the south end.

The  soft silicon gas kits are working well with the SS can.  Den is making  the adaptor plate drawing for the feedthrough.

 To put everything in one place I add a final drawing of the base to this elog.

 Next time we continue with wiring and putting temperature and pressure sensors inside the box. Connector support plate drawing is attached. We'll have sensors inside the kit with STS-2 or Trillium as their connector is small enough (19 pin vs 26 pin for Guralps) that we can put an additional 4 pin lemo connecor (2 pins for each sensor). I think EGG.0B.304.CLL is good for this application. Temperature and pressure sensor we can by from omega.

  8476   Tue Apr 23 15:02:19 2013 Max HortonUpdateSummary PagesImporting New Code

Duncan Macleod (original author of summary pages) has an updated version that I would like to import and work on.  The code and installation instructions are found below.

I am not sure where we want to host this.  I could put it in a new folder in /users/public_html/  on megatron, for example.  Duncan appears to have just included the summary page code in the pylal repository.  Should I reimport the whole repository?  I'm not sure if this will mess up other things on megatron that use pylal.  I am working on talking to Rana and Jamie to see what is best.

http://www.lsc-group.phys.uwm.edu/cgit/lalsuite/tree/pylal/bin/pylal_summary_page
https://www.lsc-group.phys.uwm.edu/daswg/docs/howto/lal-install.html
  8475   Tue Apr 23 15:00:20 2013 JenneUpdate40m Upgrading4 pins enough?

Quote:

Are 4 of these spring loaded pins enough?  I'm not sure how one pin can hold 2 lids at each point.  It seems like we need 8 pins.

 Steve has explained to me that the pins will go in between the 2 lids, with a big washer, so that one pin holds both lids at the same time.  4 is the right number.

  8474   Mon Apr 22 20:17:05 2013 CharlesUpdateISSNew Servo w/switching filters

 

In my previous post here, a new servo design was discussed. Although the exact design used will depend on the particular noise requirements for the 40m and the Bridge Labs (requirements will be considered separately for each application), I still have to yet to see those formalized. Despite this, I have been simulating an example servo circuit with three switchable stages. The design can be found at: New Servo.

Essentially, this circuit consists of three unity gain buffers that can be switched into different filtering states. Attached is a plot of the transfer function of this particular circuit with successive stages turned on. The curve (0) corresponds to all of the filters being switched off, so the total behavior is that of a unity gain buffer. The curve (1) corresponds to the first stage being turned on with the 2nd and 3rd still acting as unity gain buffers. This first state has a gain of ~80 dB at DC and a pole at ~10 Hz which sets the unity gain crossing at ~100 kHz. The curves (2) and (3) correspond to the second and third stage being turned on, respectively. Each of these stages has a pole at DC (i.e. ~infinite gain) and a zero at 10^4 Hz. For f > 10^4 Hz, these stages have gain ~ 1, as we can see in the transfer function below.

I have also performed some noise analysis of this circuit. Attached are a few plots produced by LISO showing the resistor and op-amp noise separately (it was too cluttered on one plot) at the output node of the servo. Both of these plots have a "Sum Noise" trace, which is the sum for every circuit element and is thus identical between plots. The third noise spectrum included is simply the noise at the output referenced to the input with the previously computed transfer function. I'm not sure if there is a simple method embedded in LISO to reference the noise at the output node to the input, but it should be as simple as numerically dividing the noise spectrum by the transfer function between input and output. 

Next, I will be attempting time-dependent simulations of this simple circuit using delayed switches instead of manually controlled ones.

  8473   Mon Apr 22 19:48:56 2013 JenneUpdate40m Upgrading4 pins enough?

Are 4 of these spring loaded pins enough?  I'm not sure how one pin can hold 2 lids at each point.  It seems like we need 8 pins.

  8472   Mon Apr 22 17:43:09 2013 SteveUpdate40m Upgrading ETMY optical table & enclosure is ready for optics

Quote:

Quote:

 Enclosure is at the east end. It has it's bottom o-ring in place. It will be ready for optics tomorrow around 5pm

I have to shim out the enclosure, finish leveling the table and cut surgical tubing O-ring for the top.

 

 Glued surgical latex tubing with super glue into O-ring shape. The existing in place tubing K-100, OD 0.125" (actual size 0.140"), wall 0.031", ID 0.062".

I have just found out that tolerances on tubing OD are + - 0.026" by the manufacturer. I'm getting larger tubing for better fit.

The table is ready for optics.

Things left to do:

1, finalize o-ring size  2, finish cable feedthrough  3, finalize window connection 4, IR-Thermashield strips for bridge sides 5, replace bridge support post with solid one

 

 I'm working on to improve the quality of the enclosure.

The short comings are: more cable feedthroughs needed, latches to anchor top covers air tight and posts to support bending bridges.

Red triangles are compression latches at 10 places to hold the top air tight on surgical tubing

Green lines represent 4  posts of Al 1" OD to support the covers and maximize their eigenfrequencies.

Black crosses are 4 spring loaded push-bottom quick release pins to anchor the top covers to the bridges. This connection will  not be air tight.

(quarter-turn wing head fastener have the same problem) I'm thinking of some solution to minimize the leak.    

Violet _ steel plate (1" wide, 15" long, 0.125" thick)  between the two posts will anchor the quick release pins and make bridge rigid.

Blue rectangle is an other cable feedthrough exiting on the chamber side.

Planning to substitute window with soft - air tight ( Aluminized thin wall hose )  connection to vacuum view port where white circle is representing the Al adaptor ring.

Updated after Wednesday meeting 4-24-2013

 

  8471   Mon Apr 22 17:06:42 2013 ranaSummaryIOOMC locked/aligned. MC WFS offloading by ezcaservo

 

 Why use the PSL beam as a reference? Don't we want to keep the MC pointing in a good direction through the Faraday instead???

  8470   Mon Apr 22 12:03:58 2013 KojiUpdatePSLPMC aligned too

PMC aligned. C1:PSL-PMC-PMCTRANSPD improved from 0.72ish to 0.835ish.

  8469   Mon Apr 22 11:46:09 2013 KojiSummaryIOOMC locked/aligned. MC WFS offloading by ezcaservo

Еру ьс шы тщц дщслув фтв фдшптувю

Фдыщ ш кфт еру ащддщцштп ыскшзе ещ щаадщфв еру ЬС ЦАЫ ыукмщю

I blame Den for russian keyboard installation on the control machines.

ezcaservo -r 'C1:SUS-MC2_ASCPIT_OUT16' -g '0.00001' -t 60 C1:SUS-MC2_PIT_COMM&
ezcaservo -r 'C1:SUS-MC2_ASCYAW_OUT16' -g '0.00001' -t 60 C1:SUS-MC2_YAW_COMM&
ezcaservo -r 'C1:SUS-MC1_ASCPIT_OUT16' -g '0.00001' -t 60 C1:SUS-MC1_PIT_COMM&
ezcaservo -r 'C1:SUS-MC1_ASCYAW_OUT16' -g '0.00001' -t 60 C1:SUS-MC1_YAW_COMM&
ezcaservo -r 'C1:SUS-MC3_ASCPIT_OUT16' -g '0.00001' -t 60 C1:SUS-MC3_PIT_COMM&
ezcaservo -r 'C1:SUS-MC3_ASCYAW_OUT16' -g '0.00001' -t 60 C1:SUS-MC3_YAW_COMM&

  8468   Mon Apr 22 11:26:25 2013 KojiConfigurationCDSsome RT processes restarted

When I came to the 40m, I found most of the FB signals are dead.

The suspensions were not dumped but not too much excited. Use watchdog switches to cut off the coil actuators.

Restarted mxstream from the CDS_FE_STATUS screen. The c1lsc processes got fine. But the FB indicators for c1sus, c1ioo, c1iscex/y are still red.

Sshed into c1sus/ioo, run rtcds restart all . This made them came back under control.

Same treatment for c2iscex and c1iscey. This made c1sus stall again. Also c1iscey did not come back.

At this point I decided to kill all of the rt processes on c1sus/c1ioo/c1iscex/c1iscey to avoid interference between them.
And started to restart from the end machines.

c1iscex did not come back by rtcds restart all.
Run lsmod on c1iscey and found c1x05 persisted stay on the kernel. rmmod did not remove the c1x05 module.
Run software reboot of c1iscey. => c1iscey came back online.

c1iscey did not come back by rtcds restart all.
Run software reboot of c1iscex. => c1iscex came back online.

c1ioo just came back by rtcds restart all.

c1sus did not come back by rtcds restart all.
Run software reboot of c1sus => c1sus came back online.

This series of restarting made the fb connections of some of the c1lsc processes screwed up.
Run the following restarting commands => all of the process are running with FB connection.
rtcds restart c1sup
rtcds restart c1ass
rtcds restart c1lsc

Enable damping loops by reverting the watchdog switches.

All of the FE status are green except for the c1rfm bit 2 (GE FANUC RFM CARD 0).

  8467   Fri Apr 19 16:58:59 2013 JenneUpdateASCArm A2L measurement scripts 90% working again

After Den's work with the ASS model this week, all of the channel names were changed (this wasn't pointed out in his elog....grrr), so none of the A2L scripts worked. 

They are now back, however there is still some problem with the plotting that I'm not sure I understand yet.  So, the measurement works, but I don't think we're saving the results and we certainly aren't plotting them yet. 

I wanted to check where the spots are on the mirrors, to make sure Den's stuff is doing what we think it's doing.  All of the numbers were within ~1.5mm of center, although Rossa keeps crashing (twice this afternoon?!?), so I can't copy and paste the numbers into the elog.

A near-term goal is to copy over Den's work on the Yarm to the Xarm, so that both arms will auto-align.  Also, I need to put the set of alignment scripts in a wrapper, and have that wrapper call-able from the IFO Configure screen.

Also, while thinking about the IFO Configure screen, the "save" scripts weren't working (on Rossa) today, even though I just made them work a week or so ago. Rossa, at least, was unhappy running csh, so I changed the "save" script over to bash.

  8466   Fri Apr 19 15:19:25 2013 JamieUpdatePEMTrilliums moved from bench to concrete

I moved the two Trillium seismometers that Den left on the electronics bench out onto the new concrete blocks in the lab that will be their final resting places.  I moved one onto the slab at the vertex and the other to the slab at the Y end.  I left them both locked and just sitting on the concrete.

The pile of readout electronics that were sitting next to them I moved on to the yellow foam box half way down the MC tube, between the MC tube and the X arm tube.  This is obviously not a good place to store them, but I couldn't think of a better place to put them for the moment.

  8465   Fri Apr 19 13:28:39 2013 DenUpdateASCdither alignment of yarm

I've put 4 scripts into ASS directory for YARM alignment. They should be called from !Scripts YARM button on c1ass main medm screen.

Scripts configure the servo to align the cavity and then save computed offsets. If everything goes right, no tuning of the servo is needed.

Call TRANS MON script to monitor YARM transmission, then "ON" script for aligning the cavity, then "SAVE OFFSETS" and "OFF" for turning the servo off.

ON script:

  • sets demodulation gains that I used during OL measuments
  • sets LO oscillator frequency and amplitude for each optic
  • sets demodulation phase rotation
  • sets sensing matrix
  • sets servo gains for each degree of freedom
  • sets up limits for servo outputs
  • gently increases the common gain from 0 to 1

SAVE OFFSETS script:

  • holds servo outputs
  • sets servo common gain to 0 and clears outputs
  • reads old optics DC offsets
  • computes new DC offsets
  • writes new offsets to C1:SUS-OPTIC_ANGLE_OFFSET channel
  • holds off servo outputs

OFF script:

  • sets LO amplitudes to 0
  • blocks servo outputs

Notes:

SAVE OFFSET script writes DC offsets to C1:OPTIC_ANGLE_OFFSET channel, not to _COMM channel!

LIMITS are set to 500 for cavity axis degrees of freedom and to 0.5 for input steering. Usually servo outputs is ~30% if these numbers. But if something goes wrong, check this for saturation.

DC offsets of all 8 degrees of freedom are written one by one but the whole offset of put at the same time. This works fine so far, but we might change it to ezcastep in future.

  8464   Fri Apr 19 04:20:41 2013 DenUpdateLockingPRMI on sidebands

Tonight PRMI was locked on REFL55 I&Q for PRCL and MICH with POP110I as a trigger and power normalizer.

I could see power fluctuations and beam motion on the POP camera very much the same as for carrier. The difference is that carrier stays for hours while sidebands for a few minutes.

POP110:

I&Q analog gains were set to 15 dB. Relative phase was set to 25 degrees by looking at I and Q components when the cavity goes through the resonance. Q should be 0.

pop_iq.png

REFL55:

Phase rotation was measured by exciting PRM at 20 Hz and minimizing this line at REFL55_Q. I stopped at 33 degrees.

 RIN:

I compared power fluctuations of PRCL when it was locked on carrier (POP_DC) and on sidebands (POP110_I).

rin.png

 

Time series of POP110_I during one of the locks

pop110_i.png

POP camera:

  8463   Thu Apr 18 21:12:56 2013 ManasaUpdateLockingFixed

[Den, Manasa]

TRY & TRX power measurement was redone.

TRY measures 66uW and 0.8counts on dataviewer.
TRX measures 70.4uW and 0.84counts on dataviewer.

___________________________
Detector       Power
-------------------------------------------------
QPD-Y          33uW (50%)
TRY-PD         29.8uW (49%)
Y-Camera         1%
QPD-X         35.2uW (50%)
TRX-PD        25.1uW (90%)
X-Camera    10%
____________________________

  8462   Thu Apr 18 19:54:11 2013 JenneUpdateLSCLSC whitening triggering working

I have implemented automatic triggered switching of the analog whitening (and digital dewhitening). 

The trigger is the same as the degree of freedom trigger.  On the LSC RFPD screen there is a space to enter the amount of time (in seconds) you would like to wait between receiving a trigger and actually having the whitening filter switch. 

The trigger logic is as follows: 

* For each column of the LSC input matrix (e.g. AS11 I), check if there is a non-zero element.  If there is a non-zero element (indicating that we are using that PD as the error signal for a degree of freedom), check if the corresponding DoF has been triggered.  Repeat for all columns of the matrix. 

* If either the I or the Q signal from a single PD is being used, send a trigger in the direction of the PD signal conditioning / phase rotation blocks.  (Since the whitening happens before the phase rotation, we want to have the whitening state be the same for both the I and Q signals coming from the demod boards.

* Before actually changing the whitening state, wait for the amount of time indicated on the RFPD overview screen.

* Switch the digital dewhitening.  If the digital dewhitening is on, send a bit over to the binary I/O to switch the analog whitening on.

LSC_triggers.png

LSC_SigCond.png

 

This required changing the LSC RF_PD library part so that you can send the trigger to the filter bank from outside that part..  This part is in use by all LSC models, so I'll make sure the LLO people are aware of this change before I commit it to the svn.

RF_PD_block.png

 

While I was working on the LSC model, I also put in a wait between the time that the filter module trigger is received, and when it actually switches the filter modules.  So far, this time is defined for a whole filter bank (so all filters for a given DoF still switch at the same time).  If I need to go back and make the timing individual for each filter module, I can do that.  This new EPICS variable (the WAIT) defaults to zero seconds, so the functionality will not change for anyone who uses this part.

LSC_FM_Trig.png

These changes also require 2 pieces of c-code:  {userapps}/cds/common/src/wait.c and {userapps}/isc/c1/src/inmtrxparse.c

  8461   Thu Apr 18 15:08:14 2013 SteveUpdatePEMseismometer isolation kit in place

Quote:

 We got granite bases today from the manufacturer. We plan to set them up on Wednesday, 8 am. Please note, there will be an installation mess at Xend, Yend and corner during ~4 hours. Let us know if you have any objections to do this at this particular time.

Installation locations are specified in elog 8270, scheme attached is valid except for Xend. Instrument will be installed on the place of nitrogen containers.

(  next to the wall at corner sout-east of the south end )

 The carpenter shop finished the installation of the 3 granite bases.Rapid Set Cement All high strength non-shrink grout was used.

 Compressive strength  3000 PSI at 1 hour and 9000 PSI at day 28 The janitor is still cleaning up after them at the south end.

The  soft silicon gas kits are working well with the SS can.  Den is making  the adaptor plate drawing for the feedthrough.

  8460   Thu Apr 18 02:51:52 2013 DenUpdatePSLFSS slow servo

Today Rana pointed out that our FSS slow servo is malfunctioning. It has been for a while that our laser temperature control voltage drifted from 0 to 10.

I looked at FSSSlowServo script that runs at op340m and controls the servo. Script disables the servo when MC transmission is less then FSS_LOCKEDLEVEL. But his value was set to 0.2 probably till reference cavity time.

This means that slow servo was not disabled when MC was unlocked. I changed this value to 7000.

Also I increased integral gain from 0.0350 to 0.215 such that fast control is always in the range 4.5 - 5.5

  8459   Thu Apr 18 02:24:58 2013 DenUpdateASCdither alignment of yarm

I modified our existing c1ass model to include alignment of input steering TT1 and TT2 for YARM and BS for XARM. Corresponding medm screens are also created.

Dithering:

ETM_PIT: frequency = 6 Hz, amplitude = 100 cnts
ETM_YAW: 8 Hz, 400 cnts
ITM_PIT: 11 Hz, 800 cnts
ITM_YAW: 14 Hz, 1200 cnts

These values were chosen by looking at cavity transmission and length signals - excitation peaks should be high enough but do not shake the optics too much.

Demodulation:

LO for each degree of freedom is mixed with cavity length and transmission signals that are first bandpassed at LO frequency. After mixing low-pass filter is applied. Phase rotation is chosen to minimize Q component

ETM_PIT_LENGTH 0
ETM_YAW_LENGTH 20
ITM_PIT_LENGTH 0
ITM_YAW_LENGTH -25
ETM_PIT_TRANS -5
ETM_YAW_TRANS 10
ITM_PIT_TRANS 10
ITM_YAW_TRANS -30

Sensing matrix:

8 * 8 matrix was measured by providing excitation at 0.03 Hz to optics and measuring the response in the demodulated signals. Excitation amplitude was different for each optics to create cavity transmission fluctuations of 25%

-0.0373333 -0.010202 -0.018368 0.0042552 0 0 0 0
0.0432509 -0.209207 0.0139471 0.0780632 0 0 0 0
0.0483903 -0.0077304 0.00917147 0.000860323 0 0 0 0
-0.0751211 0.699778 -0.0115889 -0.09944 0 0 0 0
0.356164 0.121226 0.0690162 -0.0183074 -59.52 -21.9863 -30.9437 13.5582
-0.141744 1.15369 -0.0100607 -0.12914 -18.8434 -105.828 -48.213 14.8612
-0.0446516 0.00682156 -0.0204571 -0.00207764 21.3057 -1.66971 22.1538 3.93419
0.0278091 -0.205367 0.0114271 0.0648548 -4.66919 97.9043 -6.26847 -95.9963

Though coherence was > 0.95 during the measurement for each element (except for TT -> Length signals), after inverting and putting it to control servo, loops started to fight each other. So I decided to try a simple diagonal matrix:

TT1_PIT -> ETM_PIT_TRANS, TT1_YAW -> ETM_YAW_TRANS, TT2_PIT -> ITM_PIT_TRANS, TT2_YAW -> ITM_YAW_TRANS,

ITM_PIT -> ETM_PIT_LENGTH, ITM_YAW -> ETM_YAW_LENGTH, ETM_PIT -> ITM_PIT_LENGTH, ETM_YAW -> ITM_YAW_LENGTH

And this matrix worked much better.

Control loops:

8 loops are running at the same time. UGF for input steering loops is 20 mHz, for cavity axis loops - 80 mHz. Slower loop is stronger at low frequencies so that cavity axis servo follows input steering alignment.

OL.png

Results:

When I started experiment the cavity was misaligned, transmission was ~0.4. Servo was able to align the cavity in ~30 seconds. This time depends on mirrors misalignment as well as input optics and cavity axis misalignment relative to each other.

When servo converged I disturbed ETMY, ITMY, TT1 and TT2. Servo was able to compensate for this.

dither_yarm.png

Excitation lines seen by transmission and length of the cavity are suppressed as shown on the attached as pdf figures.

Note:

Though the servo is able to align the cavity during my tests, this does not mean it will work perfectly any time. So please, if you lock, try to use the servo for alignment. If something goes wrong we'll fix it. This is better then to align IFO by hands every time.

  8458   Wed Apr 17 02:20:13 2013 ManasaUpdate40m UpgradingEndtable upgrade for auxiliary green laser : progress

 

Assembly progress:

1. ETMY oplev setup has been put together. Because of the shift in the TRY path, I had to modify the oplev path on the table as well.

TRY_OPL.jpg

2. Green laser setup coming together:
    (i) Used a HWP after the NPRO to convert s-polarization to p-polarization. (Verified by introducing a PBS after the HWP and then removed later).
    (ii) Lens focuses the beam into the Faraday.
    (iii) Used steering mirrors to align the beam to the faraday. With 320mW before the Faraday, I was able to get 240mW after the output aperture. The spec sheet for the faraday specifies a 93% transmission; but what I measure is only 75%.

GRY.jpg

 

  8457   Mon Apr 15 17:44:28 2013 AlbertUpdateoptical tablesOptical Table Toolboxes Update

I decided to go see what the electrical tape looks like on the other tools.

These are the tools I felt were necessary to label with tape: (the others don't seem to be terribly important in terms of not interchanging between boxes)

photo.JPG

On another note I'm not sure why electrical tape can't be used on the Allen Wrenches too.

I also plan on ordering smaller flash lights for each table (this one is bulky and unwieldy), and filling in the gaps of the Allen Wrench sets as soon as I get the go-ahead.

  8456   Mon Apr 15 16:10:52 2013 DenUpdatePEMseismometer isolation kit

 We got granite bases today from the manufacturer. We plan to set them up on Wednesday, 8 am. Please note, there will be an installation mess at Xend, Yend and corner during ~4 hours. Let us know if you have any objections to do this at this particular time.

Installation locations are specified in elog 8270, scheme attached is valid except for Xend. Instrument will be installed on the place of nitrogen containers.

(  next to the wall at corner sout-east of the south end )

  8455   Sun Apr 14 23:20:42 2013 DenUpdateLockingFixed

Quote:

TRY path fixed and ready for normalization.

I used 2" BS at R=50 and R=98 to reflect the Y arm transmission at QPD-Y and TRY PD respectively. The residual beam transmitted by the BS is now steered by a Y1mirror to the camera. With Y arm locked, transmission currently measures 40mW against the expected 70mW. TRY shows 0.45 counts in dataviewer.

 I think it is too much. Incident power to IFO is 1.3 W. Even if we assume no losses and pick-offs on the path to the arms, we should get ~100 uW out of the cavity. I measured X and Y arms transmission to be 60 uW. Did you disable triggering during your measurement?

  8454   Sun Apr 14 17:56:03 2013 ranaUpdateLockingprcl angular motion

Quote:

Quote:

For the PRM, it is also a mostly translation effect as calculated at the PRC waist position (ITM face).

I made another estimation assuming that PRCL RIN is caused by translation of the cavity axis:

  • calibrated RIN to translation, beam waist = 4mm

 In order to get translation to RIN, we need to know the offset of the input beam from the cavity axis...

This should be possible to calibrate by putting a pitch and yaw excitation lines into the PRM and measuring the RIN.

See secret document from Koji.

  8453   Sun Apr 14 17:30:14 2013 ManasaUpdateLockingFixed

TRY path fixed and ready for normalization.

I used 2" BS at R=50 and R=98 to reflect the Y arm transmission at QPD-Y and TRY PD respectively. The residual beam transmitted by the BS is now steered by a Y1mirror to the camera. With Y arm locked, transmission currently measures 40mW against the expected 70mW. TRY shows 0.45 counts in dataviewer.

  8452   Sun Apr 14 15:03:17 2013 ManasaUpdateLockingFixing - progress

Quote:

TRY signals are all gone!  Both the PD and the camera show no signal.  I went down there to turn off the lights, and look to see what was up, and I don't see any obvious things blocking the beam path on the table.  However, Steve has experimentally bungeed the lids down, so I didn't open the box to really look to see what the story is.

Absent TRY, I redid the IFO alignment.  Yarm locked, so I assumed it was close enough.  I redid Xarm alignment pretty significantly.  Transmission was ~0.5, which I got up to ~0.85 (which isn't too bad, since the PMC transmission is 0.74 instead of the usual 0.83).  I then aligned MICH, and PRM.  After fixing up the BS alignment, the POP beam wasn't hitting the POP PD in the center any more.  I centered the beam on the PD, although as Gabriele pointed out to me a week or two ago, we really need to put a lens in front of POP, since the beam is so big.  We're never getting the full beam when the cavity flashes, which is not so good.

Den is still working on locking, so I'll let him write the main locking report for the night.

We see that the PRC carrier lock seems to be more stable when we lock MICH with +1 for ITMY and -1 for ITMX, and PRCL with -1 for both ITMs.  This indicates that we need to revisit the systematic problem with using the PRM oplev to balance the coils, since that oplev has a relatively wide opening angle.  I am working on how to do this.

I'm fixing the TRY path.

I misaligned PRM and restored ETMY; but did not see the Y arm flashing. I am going ahead and moving the optics to get Y arm flashing again.

The slider values on the medm screen before touching any of them (for the record):

       tt1        tt2      itmy        etmy
p    -1.3886    0.8443     0.9320    -3.2583
y     0.3249    1.1407    -0.2849    -0.2751
  8451   Sat Apr 13 23:11:04 2013 DenUpdateLockingprcl angular motion

Quote:

For the PRM, it is also a mostly translation effect as calculated at the PRC waist position (ITM face).

I made another estimation assuming that PRCL RIN is caused by translation of the cavity axis:

  • calibrated RIN to translation, beam waist = 4mm
  • measured PRM yaw motion using oplev
  • estimated PR3 TT yaw motion: measured BS yaw spectrum with oplev OFF, divided it by pendulum TF with f0=0.9 Hz, Q=100 (BS TF), multiplied it by pendulum TF with f0 = 1.5 Hz, Q = 2 (TT TF with eddy current damping), accounted for BS local damping that reduces Q down to 10.

PRM and TT angular motion to cavity axis translation I estimated as 0.11 mm/urad and 0.22 mm/urad assuming that TTs are flat. We can make a more detailed analysis to account for curvature.

I think beam motion is caused by PR3 and PR2 TT angular motion. I guess yaw motion is larger because horizontal g-factor is closer to unity then vertical.

  8450   Sat Apr 13 03:45:51 2013 ranaUpdateLockingprcl angular motion

 

 Maybe its equivalent, but I would have assumed that the input beam is fixed and then calculate the cavity axis rotation and translation. If its small, then the modal expansion is OK. Otherwise, the overlap integral can be used.

For the ETM motion, its a purely translation effect, whereas its tilt for the ITM. For the PRM, it is also a mostly translation effect as calculated at the PRC waist position (ITM face).

  8449   Fri Apr 12 13:21:34 2013 DenUpdateLockingprcl angular motion

Quote:

 How is the cavity g-factor accounted for in this calculation?

 I assume that pointing noise and dc misalignment couples 00 to 01 by a factor theta / theta_cavity

Inside the cavity 01 is suppressed by 2/pi*F*sin(arccos(sqrt(g_cav))).

For the XARM this number is 116 taking g-factor to be 0.32. So all pointing noise couples to power RIN.

Suppression factor inside PRC is 6.5 for g-factor 0.97. This means that 85% of jitter couples to RIN, I accounted for this factor while converting RIN to angle.

I did not consider translational motion of the beam. But still PRC RIN can not be explained by oples readings as we can see exciting optics in pitch and yaw. I suspect this RIN is due to PR3, as it can create stronger motion in yaw than in pitch due to incident angle and translational motion of the mirror. I do not have a number yet.

  8448   Fri Apr 12 10:33:42 2013 CharlesSummaryISSDC-Coupled ISS Servo Design

General ISS Design

Signals through the ISS are directed as follows:  an error signal is obtained by summing the ~5 V signal from the PD with a -5 V signal from a high precision voltage regulator (which is first filtered with an ~ 30 mHz low-pass Sallen-Key filter).  It is this signal that is processed/amplified by the servo. The output from the servo is then used to drive an AOM (it is not known exactly how this is done and whether or not any preamplifier/extra circuitry is necessary). The resulting modulation, hopefully, reduces fluctuations in the laser intensity incident on the PD, lowering the relative intensity noise.

Servo Design

Almost the entirety of my focus has been directed toward designing the servo portion of the ISS. Speaking in general terms, the currently proposed design consists of stages of active op-amp filters, but now the stages will have internal switches that allow them to switch between ‘flat’ gain buffers and more complicated filters with our desired behavior. Consider some Example Filter Stages where I have demonstrated a typical switching filter with the switch open and closed. When the switch is closed, the capacitor is shorted and we simply have a variable gain buffer (variable in the sense that its gain can be tuned by proper choice of the resistances) with no frequency dependence. When the switch is open, the capacitor introduces a pole at ~100 Hz and a zero at ~1 kHz.

CircuitLab has decent analysis capabilities and attached are plots generated by CircuitLab. The first plot corresponds to a frequency analysis of the voltage gain of op-amp U1 and the ‘flat’ ~20 dBV gain filter with the switch closed and the capacitor shorted. The second plot is the same frequency analysis, but now with op-amp U2 and the filter with the switch open and the capacitor introduced into signal processing. This particular combination of resistors and capacitors produce a DC gain of 60 dBV, a pole at ~100 Hz, a zero at ~10 kHz and high frequency behavior of ~constant gain of 20 dBV. In this simulation, the gain-bandwidth product of the simulated op-amp (the standard op-amp CircuitLab uses) was artificially increased in order to see more ideal behavior in the higher frequency domain.

Switches like the above can be used to add boosts to some initial filter state (which could be like the above or possibly a simple integrator to achieve high DC gain) and change it into a more complex and more useful filter state advantageous for desired noise suppression. Cascades of these switching filters could be used to create very complicated transfer function behavior. No general servo has yet been designed as the exact details of the intensity noise requirements are still being determined.

With regards to the implementation of the switches, some ‘smart’ signal will be used to trigger a switch opening and the boost being introduced to the signal processing. The switches will be opened (open corresponds to adding the boost) in a manner that maintains stability of the servo circuit. Essentially, some sort of time delay or power monitor induced signal (power from the PD output) will be used to modify the servo's behavior.

AOM

How exactly the signal will drive the AOM for correct noise suppression is unknown currently.

 

  8447   Fri Apr 12 09:20:32 2013 ranaUpdateLockingprcl angular motion

 How is the cavity g-factor accounted for in this calculation?

  8446   Fri Apr 12 02:56:34 2013 DenUpdateLockingprcl angular motion

I compared PCRL and XARM angular motions by misaligning the cavities and measuring power RIN. Divergence angles for both cavities I calculated to be 100 urad.

XARM pointing noise sums from input steering TTs, PR2 and PR3 TTs, BS, ITMX, ETMY.

PRCL noise - from input TT, PRM, PR2 and PR3 TT, BS, ITMX, ITMY.

I would expect these noises to be the same as angular motion of different optics measured by oplves is simular. We do not have oplves on TT but they are present in both passes.

I measured RIN and converted to angle. Sharp 1 Hz resonance at XARM pointing spectrum is due to EMTX, it is not seen by PRCL. Other then that XARM is much quiter, especially at 3 - 30 Hz.

As PRM  is the main difference in two passes, I checked its spectrum. When PRCL was locked I excited PRM in pitch and yaw. I could see this excitation at RIN only when the peak was 100 times higher then background seismic noise measured by oplev.

pointing.png

  8445   Thu Apr 11 16:29:23 2013 AlbertUpdateoptical tablesOptical Table Toolboxes Update

Quote:

 There are some tips for how to appy nail polish on YouTube from MKNails and MissJenFABULOUS. Their tips on how to prepare the site for a strong bonding strength are probably helpful for our gold/nickel coated tools. For chrome tools we may need to abrade the surface with a stone or fine sandpaper for it to take the layer better. IF the YouTube videos don't do it for you, then I suggest contacting Tom Evans at LLO to find out what kind of nail polish he uses.

 This is the tentative box placement per optical table. The toolboxes are going to be color-coded by a combination of two colors (the order won't matter). The side of each toolbox will have a little panel to let you know which box corresponds to which set of colors.

On the diagram, the set of colors is simply the color of the box border and the color of the text.

If anyone has a problem with any of the colors or the box placement let me know before they are installed and become an annoyance:

Diagram.png

Box Placements:

ETMY: Box will be attached to the underside of the table by magnets. The box will be on the north side of the optical table.

POY: Box will be attached to the side of the optical table by magnets. The box will be on the west side of the optical table.

BSPRM: Box will be attached to the side of the optical table by magnets. The box will be on the west side of the optical table.

AS: Box will be attached to the side of the optical table by magnets. The box will be on the north side of the optical table.

PSL1: Box will be inside the optical table, in the northeast corner.

PSL2: Box will be inside the optical table, in the southwest corner.

POX: Box will be attached to the side of the optical table by magnets. The box will be on the south side of the optical table.

MC2: Box will be attached to the side of the optical table by magnets. The box will be on the south side of the optical table.

ETMX: Box will be attached to the side of the optical table by magnets. The box will be on the east side of the optical table.

  8444   Thu Apr 11 11:58:21 2013 JenneUpdateComputersLSC whitening c-code ready

The big hold-up with getting the LSC whitening triggering ready has been a problem with running the c-code on the front end models.  That problem has now been solved (Thanks Alex!), so I can move forward.

The background:

We want the RFPD whitening filters to be OFF while in acquisition mode, but after we lock, we want to turn the analog whitening (and the digital compensation) ON.  The difference between this and the other DoF and filter module triggers is that we must parse the input matrix to see which PD is being used for locking at that time.  It is the c-code that parses this matrix that has been causing trouble.  I have been testing this code on the c1tst.mdl, which runs on the Y-end computer.  Every time I tried to compile and run the c1tst model, the entire Y-end computer would crash.

The solution:

Alex came over to look at things with Jamie and me.  In the 2.5 version of the RCG (which we are still using), there is an optimization flag "-O3" in the make file.  This optimization, while it can make models run a little faster, has been known in the past to cause problems.  Here at the 40m, our make files had an if-statement, so that the c1pem model would compile using the "-O" optimization flag instead, so clearly we had seen the problem here before, probably when Masha was here and running the neural network code on the pem model.  In the RCG 2.6 release, all models are compiled using the "-O" flag.  We tried compiling the c1tst model with this "-O" optimization, and the model started and the computer is just fine.  This solved the problem.

Since we are going to upgrade to RCG 2.6 in the near-ish future anyway, Alex changed our make files so that all models will now compile with the "-O" flagWe should monitor other models when we recompile them, to make sure none of them start running long with the different optimization. 

The future:

Implement LSC whitening triggering!

  8443   Thu Apr 11 10:15:55 2013 SteveUpdateLockingPRM yaw oplev transferfunction

See   Feb 2012 PRM yaw transferfunctions, also check Valera's modified  side sensor may effect yaw motion

 

ELOG V3.1.3-